Ask
your own question, for FREE!
Collaborative Statistics
12 Online
Suppose that 50 identical batteries are being tested. After 8 hours of continuous use, assume that a given battery is still operating with a probability of 0.70 and has failed with a probability of 0.30. What is the probability that between 25 and 30 batteries (inclusive) will last at least 8 hours
Still Need Help?
Join the QuestionCove community and study together with friends!
Given: Bernoulli trials, (population) success rate = p = 0.7 (constant throughout trials) Assume independence of the 50 batteries use binomial distribution with parameters (n,p)=(50,0.7) \(P(X\in {25,26,27,28,29,30}) =\sum_{i=25}^{30}\left(\begin{matrix}50 \\ i\end{matrix}\right)0.7^i 0.3^{50-i}\) =0.001436+0.003223+0.006684+0.012811+0.022677 =0.037039
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
1 day ago
1 Reply
0 Medals
6 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
3 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal